You want to design an oval racetrack such that 3200 lb racecars can round the turns of radius 1000 ft at 105 mi/h without the aid of friction. You estimate that when elements like downforce and grip in the tires are considered, the cars will round the turns at a maximum of 175 mi/h . Find the banking angle ? necessary for the race cars to navigate these turns at 105 mi/h and without the aid of friction. This banking and radius are very close to the actual turn data at Daytona International Speedway where 3200 lb stock cars travel around the turns at about 175 mi/h. What additional radial force is necessary to hold the race car on the track at 175 mi/h?
Answer #1

Ans

The racecars can travel around the of radius (R)= 1000 ft = 1000*0.3048 m = 304.8 m

When there is no friction, the banking angle ? is given by

v = speed of the car = 105 mi/h = 105*0.44704 = 46.95 m/s

so the angle ? is

------------------------------------------------------------

The static friction is the additional force necessary to hold the race car.